Integrals Giving Inverse Trigonometric Functions


1. ∫(x+1)/(x² +1) from 0 to 1


2. ∫(dz)/z(1 + ln² z) from 1 to e


1
Expert's answer
2021-04-12T18:05:51-0400

1) "\int_{0}^1 \frac{x+1}{x^2 +1}dx" = "\int_{0}^1 \frac{x}{x^2 +1} + \frac{1}{x^2 + 1}dx"

= "\int_{0}^1 \frac{x}{x^2 +1}dx + \int_{0}^1\frac{1}{x^2 + 1}dx"

= "\int_{0}^1 \frac{x}{x^2 +1}dx" + tan-1(x) "|_0^1"


let x2 + 1 = u "\implies" 2x"dx" = "du" & x = 0 then u = 1

"\implies" x"dx" = "\frac{du}{2}" & x = 1 then u = 2

therefore given integral becomes,

= "\int_{1}^2 \frac{1}{2u}du" + tan-1(1) - tan-1(0)

= "\int_{1}^2 \frac{1}{2u}du" + "\frac{\pi}{4}" - 0


= "\frac{1}{2}" ln(u) "|_1^2" + "\frac{\pi}{4}"

= "\frac{1}{2}" ( ln(2) - ln(1) ) + "\frac{\pi}{4}"

= "\frac{1}{2}" ln(2) + "\frac{\pi}{4}"

2) "\int_{1}^e \frac{1}{z(1 + ln^2 (z) }dz" "\implies" let "ln(z)" = t & z = e then t = 1

"\implies" "\frac{1}{z} dz = dt" & z = 1 then t = 0


therefore integral becomes,

= "\int_{0}^1 \frac{1}{(1 + t^2) }dt"

= tan-1(t) "|_0^1\u2223 \n\n\u200b"


= tan-1(1) - tan-1(0)


= "\frac{\pi}{4}"